OpenStudy (anonymous):

Factor the following polynomial completely. -x 2 y 2 + x 4 + 9y 2 - 9x 2

OpenStudy (anonymous):

(1) pq - 3r + pr - 3q = (pq + pr) - 3(r + q) = p(q + r) - 3(q + r) = (q + r)(p - 3) (2) -x^2y^2 + x^4 + 9y^2 - 9x^2 = -x^2(y^2 - x^2) + 9(y^2 - x^2) = (y^2 - x^2)(9 - x^2) = (y + x)(y - x)(3 + x)(3 - x) got it from yahoo answers